# Bilateral Distribution Compression: Reducing Both Data Size and Dimensionality

### v4

@@ 4.3 Exact Bilateral Distribution Compression
In previous experiments, the EMMD, RMMD, and DMMD must be estimated, but for certain kernel–distribution pairs analytical formulae are available (Briol et al., 2025). To demonstrate that BDC effectively minimises the DMMD, we generate n=20,000n=20{,}000 samples from a two-dimensional mixture of Gaussians and project them to $d=250$ dimensions using a matrix $V\in\mathbb{R}^{250\times 2}$ of standard Gaussian draws. With a Gaussian kernel, the EMMD, RMMD, and DMMD can then be evaluated exactly (see Section D.2). We apply BDC to this 250-dimensional dataset using a linear autoencoder with intrinsic dimension $p=2$; results are shown in Figure 11. The figure shows that the BDC-compressed set, when decoded into ambient space, achieves MMD which is equal to that achieved by the full dataset, while requiring only $0.018\%$ of the storage, at the largest size. We also empirically verify the bound from Theorem 3.5 as the pull-back kernel is used here.
Figure 11: BDC on a Gaussian mixture dataset. Left: ambient space MMD achieved by BDC (red) and ADC (blue), with the RMMD++EMMD bound (orange), each reported over 5 runs. Compared with URS over 250 runs (shaded bands: 25th–75th percentiles), with FULL as baseline (black). Right: Full dataset projected into latent space (black), with BDC compressed set of size 10 (red).
@@ 5 Ablation Studies
The algorithms developed in this work involve several hyperparameters that require selection and tuning. Due to space constraints, we report extensive ablations in the appendix, covering the choice of kernel functions (Section D.1.2), the latent dimension $p$ (Sections D.1.5, D.1.6, and D.1.7), the compressed set size $m$ (Section D.1.8), and a weighting parameter in the neural autoencoder hybrid loss (Section D.1.9). These results demonstrate that BDC is robust to sensible hyperparameter choices. We also provide practical guidance on hyperparameter selection in Section C.4. We also include a comparative study showing that BDC substantially outperforms a naive sequential baseline that applies ADC followed by PCA, both in test performance and computational efficiency (Section D.1.10).
@@ 6 Conclusions
We introduced bilateral distribution compression, which seeks to reduce both the sample size and feature dimensionality of a dataset while targeting distributional preservation. We proposed a tractable linear-time two-stage framework targeting the DMMD. The first stage learns a low-dimensional representation by minimising the RMMD, and the second stage selects a compressed set in latent space by minimising the EMMD. We established theoretical guarantees linking these quantities to the DMMD, and demonstrated empirically that the resulting compressed sets achieve comparable or superior performance across a range of tasks versus ambient-space distribution compression methods, with much reduced runtime. A detailed discussion of future work, limitations, and related work is given in Section A.
@@ Acknowledgements
We have taken several measures to ensure the reproducibility of our results. Complete algorithmic details for BDC are provided in Section 3.1, with pseudocode in Algorithms 1 and 2. Theoretical guarantees, assumptions, and proofs are given in the appendix. All datasets, preprocessing steps, and experimental setups are described in detail in Section 4 and D.1. A code repository containing an implementation of BDC is provided in the supplementary material.
@@ A.1.1 Standard Distribution Compression
Kernel Thinning (Dwivedi and Mackey, 2021a, b; Shetty et al., 2022) constructs a compressed set that is a proper subset of the original dataset, i.e., all points in the compressed set are drawn directly from the input data. This restriction stems from the method’s original motivation, thinning the output of Markov Chain Monte Carlo (MCMC) methods, where subsampling is commonly referred to as standard thinning. Kernel Thinning proceeds via a two-stage procedure targeting the MMD. For a user-specified choice of $\alpha\in\mathbb{N}$, firstly the input dataset is probabilistically split into $2^{\alpha}$ MMD-balanced candidate sets, each of size $\lfloor n/2^{\alpha}\rfloor$, discarding the remaining $n-2^{\alpha}\lfloor n/2^{\alpha}\rfloor$ points if $n$ is not evenly divisible. Secondly, the best candidate set from this partitioning is selected, and then greedily refined by iteratively replacing points with others from the original dataset whenever doing so improves the MMD. This two-stage construction admits convergence rates that are state of the art (Dwivedi and Mackey, 2021b), however, it is constrained to producing compressed sets of size $\lfloor n/2^{\alpha}\rfloor$.
@@ A.1.2 Dataset Condensation
This method exploits the latent space of randomly initialised CNNs to optimise a compressed set of synthetic images in the original input space. Empirically, training a CNN on such a compressed set achieves strong classification accuracy (Zhao and Bilen, 2022). It is important to note, however, that the objective in (6) is not an MMD: it compares only the first moment of the embedded distributions, and therefore does not measure distributional equality. A subsequent refinement of this approach by Zhao et al. (2023) incorporated augmentation, upsampling, and regularisation techniques, though the underlying methodology remained unchanged.
@@ A.2.4 General Feature Domains
In this work we have restricted our study of bilateral distribution compression to the setting where data lie in a real ambient space $\mathbb{R}^{d}$. While the framework naturally extends to more general feature domains $\mathcal{X}$, provided a positive definite kernel can be defined, we have not explored such cases empirically. In principle, the use of an autoencoder may allow us to extend gradient-based distribution compression to domains where direct gradient computation on the raw data is difficult or infeasible, such as graphs, text, and other structured or discrete inputs, where non-gradient methods like Kernel Thinning (Dwivedi and Mackey, 2021b) would be the standard. In these settings, the autoencoder maps the data into a continuous latent space where gradient-based optimisation is tractable, potentially enabling efficient, high-quality compression even in challenging modalities. We leave a detailed investigation of this potential capability to future work.
@@ B.1 Technical Assumptions and Additional Notation
$\left\|\bm{X}^{\top}\bm{X}-VV^{\top}\bm{X}^{\top}\bm{X}VV^{\top}\right\|_{F}^{2}$ $=\|\bm{XX}^{\top}\|_{2}^{2}-2\langle\bm{X}^{\top}\bm{X},VV^{\top}\bm{X}^{\top}\bm{X}VV^{\top}\rangle_{F}$
@@ 4. Step 4:
We have presented a standard gradient descent approach, but this can be extended to adaptive optimisers such as ADAM (Kingma and Ba, 2017) (see, e.g., Bécigneul2019Riemannian; Li et al. (2020)). In practice, we use vanilla ADAM on gradients projected onto the Stiefel tangent space at each step, followed by a QR retraction to restore the orthonormality constraint. This provides a practical approximation to a full Stiefel-constrained ADAM, which also applies vector transport to momentum and second-moment estimates. In our setting, the simpler projected-gradient variant performs well while being much easier to implement.
@@ C.6 Pseudocode
   Project $G$ onto the tangent space of the Stiefel manifold at $W$
   Update $V\leftarrow\mathrm{Retr}(V-\eta G)$, where $\mathrm{Retr}$ is a retraction onto the Stiefel manifold
@@ Appendix D Experiment Details
For all experiments, we use the Adam optimiser (Kingma and Ba, 2017) as a sensible default choice. However, the implementations allow for an arbitrary choice of optimiser via the Optax package (DeepMind et al., 2020).
@@ D.1.2 Additional Regression Experiment: Swiss-Roll with Linear Projection and IMQ Kernel
In this section, we revisit the linear Swiss-Roll dataset from Section D.1.1, this time using the inverse multi-quadratic (IMQ) kernel instead of the Gaussian kernel for distribution compression. IMQ kernels are characteristic (Weinstein and Miller, 2021) and take the form:

### v5

@@ 4.3 Exact Bilateral Distribution Compression
In previous experiments, the EMMD, RMMD, and DMMD must be estimated, but for certain kernel–distribution pairs analytical formulae are available (Briol et al., 2025). To demonstrate that BDC effectively minimises the DMMD to the true distribution $\mathbb{P}_{X}$, we use a Gaussian kernel, and generate n=20,000n=20{,}000 samples from a two-dimensional mixture of Gaussians, projecting them to $d=250$ dimensions using a matrix $V\in\mathbb{R}^{250\times 2}$ of standard Gaussian draws. In this case our MMDs can be evaluated exactly (see Section D.2). We apply BDC to this 250-dimensional dataset using a linear autoencoder with intrinsic dimension $p=2$; results are shown in Figure 11. The figure shows that the BDC-compressed set, when decoded into ambient space, achieves exact MMD which is equal to that achieved by the full dataset, while requiring only $0.018\%$ of the storage. We also empirically verify the bound from Theorem 3.5 as the pull-back kernel is used here.
Figure 11: BDC on a Gaussian mixture dataset. Left: ambient space MMD achieved by BDC post-decoding (red) and ADC (blue), with the RMMD++EMMD bound (orange), each reported over 5 runs. Compared with URS over 250 runs (shaded bands: 25th–75th percentiles), with FULL as baseline (black). Right: Full dataset projected into latent space (black), with BDC compressed set of size 10 (red).
@@ 5 Ablation Studies
The algorithms developed in this work involve several hyperparameters that require selection and tuning. Owing to space constraints, we present extensive ablation studies in the appendix, examining the choice of kernel functions (Section D.1.2), the latent dimension $p$ (Sections D.1.5, D.1.6, and D.1.7), and the compressed set size $m$ (Section D.1.8). In Section  D.1.9, we also study a weighting parameter in the neural hybrid loss (Equation 5). These studies demonstrate that BDC is robust to sensible hyperparameter choices. Practical guidance on hyperparameter selection is provided in Section C.4. Finally, we include a study showing that BDC substantially outperforms a naive sequential baseline that applies ADC followed by PCA, both in terms of test performance and computational efficiency (Section D.1.10).
@@ 6 Conclusions
We introduced bilateral distribution compression, which seeks to reduce both the sample size and feature dimensionality of a dataset while targeting distributional preservation. We proposed a linear-time two-stage framework targeting the DMMD. The first stage learns a low-dimensional representation by minimising the RMMD, and the second stage optimises a compressed set in latent space by minimising the EMMD. We established theoretical guarantees linking these quantities to the DMMD, and demonstrated empirically that the bilaterally compressed sets achieve comparable or superior performance across a range of tasks versus ambient-space distribution compression methods, with much reduced runtime. A detailed discussion of future work, limitations, and related work is given in Section A.
@@ Acknowledgements
We have taken several measures to ensure the reproducibility of our results. Complete algorithmic details for BDC are provided in Section 3.1, with pseudocode in Algorithms 1 and 2. Theoretical guarantees, assumptions, and proofs are given in the appendix. All datasets, preprocessing steps, and experimental setups are described in detail in Section 4 and D.1. A code repository containing an implementation of BDC is provided at https://github.com/bilateraldummy/bilateral_dummy.
Appendix Overview
  • Appendix A: Further discussion, including related work, limitations, and future directions.
  • Appendix B: Proofs of all theoretical results, together with a discussion of the restrictiveness of the technical assumptions.
  • Appendix C: Additional algorithmic details, including orthonormal manifold optimisation, the extension of BDC to labelled data, hyperparameter selection guidelines, a complexity analysis, and full pseudocode.
  • Appendix D: Full details of the experimental setup, additional experiments on further datasets, and extensive ablation studies covering kernel choice, latent dimensionality, and compressed set size.
@@ A.1.1 Standard Distribution Compression
Kernel Thinning (Dwivedi and Mackey, 2022, 2021; Shetty et al., 2022) constructs a compressed set that is a proper subset of the original dataset, i.e., all points in the compressed set are drawn directly from the input data. This restriction stems from the method’s original motivation, thinning the output of Markov Chain Monte Carlo (MCMC) methods, where subsampling is commonly referred to as standard thinning. Kernel Thinning proceeds via a two-stage procedure targeting the MMD. For a user-specified choice of $\alpha\in\mathbb{N}$, firstly the input dataset is probabilistically split into $2^{\alpha}$ MMD-balanced candidate sets, each of size $\lfloor n/2^{\alpha}\rfloor$, discarding the remaining $n-2^{\alpha}\lfloor n/2^{\alpha}\rfloor$ points if $n$ is not evenly divisible. Secondly, the best candidate set from this partitioning is selected, and then greedily refined by iteratively replacing points with others from the original dataset whenever doing so improves the MMD. This two-stage construction admits convergence rates that are state of the art (Dwivedi and Mackey, 2021), however, it is constrained to producing compressed sets of size $\lfloor n/2^{\alpha}\rfloor$.
@@ A.1.2 Dataset Condensation
This method exploits the latent space of randomly initialised CNNs to optimise a compressed set of synthetic images in the original input space. Empirically, training a CNN on such a compressed set achieves strong classification accuracy (Zhao and Bilen, 2023). It is important to note, however, that the objective in (6) is not an MMD: it compares only the first moment of the embedded distributions, and therefore does not measure distributional equality. A subsequent refinement of this approach by Zhao et al. (2023) incorporated augmentation, upsampling, and regularisation techniques, though the underlying methodology remained unchanged.
@@ A.2.4 General Feature Domains
In this work we have restricted our study of bilateral distribution compression to the setting where data lie in a real ambient space $\mathbb{R}^{d}$. While the framework naturally extends to more general feature domains $\mathcal{X}$, provided a positive definite kernel can be defined, we have not explored such cases empirically. In principle, the use of an autoencoder may allow us to extend gradient-based distribution compression to domains where direct gradient computation on the raw data is difficult or infeasible, such as graphs, text, and other structured or discrete inputs, where non-gradient methods like Kernel Thinning (Dwivedi and Mackey, 2021) would be the standard. In these settings, the autoencoder maps the data into a continuous latent space where gradient-based optimisation is tractable, potentially enabling efficient, high-quality compression even in challenging modalities. We leave a detailed investigation of this potential capability to future work.
@@ B.1 Technical Assumptions and Additional Notation
$\left\|\bm{X}^{\top}\bm{X}-VV^{\top}\bm{X}^{\top}\bm{X}VV^{\top}\right\|_{F}^{2}$ $=\|\bm{XX}^{\top}\|_{F}^{2}-2\langle\bm{X}^{\top}\bm{X},VV^{\top}\bm{X}^{\top}\bm{X}VV^{\top}\rangle_{F}$
@@ 4. Step 4:
We have presented a standard gradient descent approach, but this can be extended to adaptive optimisers such as ADAM (Kingma and Ba, 2015) (see, e.g., Bécigneul2019Riemannian; Li et al. (2020)). In practice, we use vanilla ADAM on gradients projected onto the Stiefel tangent space at each step, followed by a QR retraction to restore the orthonormality constraint. This provides a practical approximation to a full Stiefel-constrained ADAM, which also applies vector transport to momentum and second-moment estimates. In our setting, the simpler projected-gradient variant performs well while being much easier to implement.
@@ C.6 Pseudocode
   Project $G$ onto the tangent space of the Stiefel manifold at $V$, denote it as $\mathcal{G}_{V}$
   Update $V\leftarrow\mathrm{Retr}(V-\eta\mathcal{G}_{V})$, where $\mathrm{Retr}$ is a retraction onto the Stiefel manifold
@@ Appendix D Experiment Details
For all experiments, we use the Adam optimiser (Kingma and Ba, 2015) as a sensible default choice. However, the implementations allow for an arbitrary choice of optimiser via the Optax package (DeepMind et al., 2020).
@@
Wherever a Gaussian Process is applied for downstream inference, we use the GaussianProcessRegressor or GaussianProcessClassifier class from the package scikit-learn (Pedregosa et al., 2011), choosing a Gaussian kernel.
@@ D.1.2 Additional Regression Experiment: Swiss-Roll with Linear Projection and IMQ Kernel
In this section, we revisit the linear Swiss-Roll dataset from Section D.1.1, this time using the inverse multi-quadratic (IMQ) kernel instead of the Gaussian kernel for distribution compression. IMQ kernels are characteristic (Weinstein and Miller, 2023) and take the form: